# Learning-Based Difficulty Calibration for Enhanced Membership Inference Attacks

### v1

@@ Abstract
Machine learning models, in particular deep neural networks, are currently an integral part of various applications, from healthcare to finance. However, using sensitive data to train these models raises concerns about privacy and security. One method that has emerged to verify if the trained models are privacy-preserving is Membership Inference Attacks (MIA), which allows adversaries to determine whether a specific data point was part of a model’s training dataset. While a series of MIAs have been proposed in the literature, only a few can achieve high True Positive Rates (TPR) in the low False Positive Rate (FPR) region ($0.01\%\sim 1\%$). This is a crucial factor to consider for an MIA to be practically useful in real-world settings. In this paper, we present a novel approach to MIA that is aimed at significantly improving TPR at low FPRs. Our method, named learning-based difficulty calibration for MIA (LDC-MIA), characterizes data records by their hardness levels using a neural network classifier to determine membership. The experiment results show that LDC-MIA can improve TPR at low FPR by up to 4x compared to the other difficulty calibration based MIAs. It also has the highest Area Under ROC curve (AUC) across all datasets. Our method’s cost is comparable with most of the existing MIAs, but is orders of magnitude more efficient than one of the state-of-the-art methods, LiRA, while achieving similar performance.
@@ 1 Introduction
Figure 1: The distribution of target samples
@@
To tackle this issue, Watson et al. proposed the use of a reference model [37], which is an additional model trained on data drawn from the same distribution as those in the training dataset of the target model. Then, each target data record is fed into both the target and reference models to obtain a loss value, respectively. By calculating the difference between the two loss values, the reference model helps calibrate the target model’s behavior on a data record. This approach is an example of a difficulty calibration-based attack, which is currently one of the most advanced MIAs. To further improve the attack performance, Carlini et al. designed a Likelihood Ratio Attack (LiRA) [3]. LiRA utilizes multiple shadow models to estimate the distribution of loss values on a target data record for models that are either trained or not trained on this data sample. Although LiRA achieves high True Positive Rates (TPRs) at low False Positive Rates (FPRs), it requires significant computation.
Our proposed attack [^1] The implementation of LDC-MIA is available at https://github.com/horanshi/LDC-MIA is motivated by these works and aims to achieve high TPRs at low FPRs while minimizing the cost for attackers. We consider two types of costs in our attack: the training cost and the data cost. The training cost is mainly dependent on the number of attacking models and their complexities, while the data cost involves the amount of data needed to train the attacking models. We strive to minimize both in our design. In our proposed attack, we also use a reference model. However, unlike other methods, we take into account the intrinsic characteristics of the target data records when calibrating the difficulty of a target data record. Furthermore, we argue that the hardness levels of data records are not universal across different classes because of their distinct data distributions. To illustrate this, we provide an example in Figure 1. The figure shows data records belonging to different classes, Airplane and Cat, the data records are from CIFAR-10 dataset[18]. Each record is represented by a marker that indicates its membership type, while its calibrated membership score is shown on the y-axis, and its membership score on the target model is shown on the x-axis. The calibrated membership score is calculated by taking the difference between a data record’s membership score on the target model and the reference model, as proposed by Watson et al. [37].
@@
Based on the figure, we can make a few important observations. First, the calibrated membership score increases the gap between the easy-to-predict non-members and hard-to-predict members. However, it also reduces the gap between easy-to-predict non-members and hard-to-predict non-members since both groups may have low calibrated membership scores. However, the membership scores along the $x-$axis can help easily differentiate between these two groups. Second, it is evident that the two classes have different optimal thresholds for the calibrated membership scores. This indicates that the hardness levels of data records are not universal across different classes. Therefore, it is necessary to adopt a more intelligent approach to determining the threshold values. Third, the data distribution also plays an important role in determining how hard it is for a data record to be classified, in addition to the intrinsic characteristics of data records. As revealed by Long et al. [24], the more neighbors a data record has in the training dataset, the easier it is for it to be correctly classified. This also means that the data record is more likely to be determined as a member by attackers.
@@
Based on these observations, we propose developing a classifier that can learn to calibrate difficulty based on the membership score on the target model, the calibrated membership score, the label of the target data record, and its neighborhood information. To train this classifier, we can use a shadow target model and a reference model trained with data records that share the same distribution as those belonging to the target model training dataset. The shadow target model would mimic the target model’s behavior in classifying members and non-members. We call the proposed attack LDC-MIA. The main contributions of this paper are threefold. (1) The proposed attack significantly improves the TPR at low FPR while minimizing the cost for attackers. We only require one shadow model and one reference model to improve the TPR. Additionally, the classifier we build is a simple model that has three fully connected layers. (2) We conduct a comprehensive characterization of the data records’ hardness levels and use these characters to train a neural network for determining membership. This learning-based calibration approach can be easily extended to integrate other features without requiring significant retraining efforts. (3) Through extensive evaluations, we provide insights into the contributions each of the characters makes to the success of our proposed attack.
@@
We conduct extensive experiments to evaluate the performance of LDC-MIA on various datasets. Specifically, we measure the TPR at low FPRs ranging from $0.01\%$ to $1\%$. This metric helps us evaluate the model’s ability to correctly identify positive instances while minimizing the number of false positive predictions. Our results show that our proposed attack achieves the highest TPR across all datasets compared to state-of-the-art MIAs, with an improvement of up to 4x. In addition, we measure the precision-recall curve to analyze how well the model performs across different recall levels while maintaining high precision. The results indicate that LDC-MIA consistently produces the highest precision values for different recall values across all datasets. For instance, LDC-MIA identifies $52.72\%$ of the members with a precision of $80\%$, which is significantly higher than what other MIAs can achieve.
@@ 2.2 Membership Inference Attacks
Adversary’s Knowledge. To align with the setting of real-world scenarios, most of MIAs[32, 39, 17, 14] are conducted under a black-box setting, where attackers only have access to the target model’s outputs. In addition to this, attackers also possess an auxiliary dataset that shares the same data distribution as the training dataset of the target model. In some MIAs[32, 30], adversaries can exploit auxiliary datasets to train shadow models or reference models to facilitate the attack.
However, there are also MIAs that target gray-box or white-box attack scenarios[27, 26, 21, 11]. In these contexts, in addition to the model’s structure and auxiliary datasets, adversaries may have knowledge of the target model’s architecture, training algorithm, and potentially other training-related details.
@@ 3.2 Design intuition
Recent state-of-the-art attacks [3, 37] have explored the difficulty level of each data record and applied parametric calibration to improve the attack performance in the low FPR region. These attacks are similar in design to our proposed method in that each target data record is individually considered when performing attacks. However, these works provided limited discussions on the impact of calibration on different types of data records with varying intrinsic properties. Motivated by this, we categorize member and non-member data records into five categories based on their difficulty levels to predict and calibrate, as shown in Figure 3: hard-to-predict member/non-member, easy-to-predict member/non-member, and hard-to-calibrate non-member.
This figure compares membership scores on the target model and their calibrated counterparts for each data record. The $x$-axis represents the membership score on the target model, while the $y$-axis shows the calibrated membership score. The calibrated membership score is calculated as the difference between the membership score on the target model and that of a reference model. Data records are also labeled with their “hardness” levels in different areas, and members and non-members are marked with different colors. In general, higher scores on the target model’s membership suggest that a data record is a member of the model’s training dataset. This idea has been used in traditional MIAs. However, by correlating these scores with the calibrated membership score, we can gain more insights into the difficulty level of a data record. It is important to note that in our discussions, the difficulty or “hardness” level of a data record is mainly determined by the model in which it is not included as part of the training dataset. Specifically, with regards to member data records, we refer to them as being hard-to-predict when they have low membership scores on the reference model, whereas they are considered easy-to-predict members when their scores are higher. Similarly, for non-members, the difficulty level is determined by their membership scores on the target model. By analyzing this figure, we can get the following insights.
@@
Even though FPR can be reduced by carefully selecting a threshold for the calibrated scores, it is difficult to eliminate all of them. In Figure 3, it can be seen that many non-members overlap with members along the $y$-axis. This can be seen more clearly in Figure 2, which shows the distribution of calibrated membership scores for both members and non-members. In this figure, the calibrated membership scores are on the right $y$-axis and share the same values with that of Figure 3. The $x$-axis shows the number of members and non-members with corresponding calibrated scores. The figure highlights that many members have low calibrated scores, making them easy-to-predict members. These members overlap with non-members near the line where the calibrated score equals 0. Within this region, the easy-to-predict members overlap with both hard-to-predict and easy-to-predict non-members, as both groups have similar outputs on the target and the reference models.
@@
Two things are learned from these observations. Firstly, attackers are likely to encounter easy-to-predict members in real-world attacks, so identifying these samples significantly improves TPR. Secondly, existing difficulty calibration based MIAs may fail to isolate such members from non-members by only considering the calibrated membership scores. But we can see from Figure 3 that moving the vertical dashed line and the horizontal one simultaneously makes it easier to differentiate between members and non-members. This indicates that the membership scores on the target model are still useful in addition to the calibrated membership scores. Based on this observation, we utilize membership scores on the target model in LDC-MIA to exclude the hard-to-predict non-members. This approach not only helps identify hard-to-predict members but also easy-to-predict members, thus improving TPR in all FPR regions.
@@
Figure 3: Based on the calibrated membership scores and membership scores on the target model, we classify target samples into five categories: hard-to-predict member/non-member, easy-to-predict member/non-member, and hard-to-calibrate non-member.
Neighbor information is also important. In Figure 3, it is evident that some non-members have high calibrated membership scores, indicating that they perform significantly better on the target model than on the reference model. These non-members are not easy-to-predict, as they have low membership scores on the reference model, nor hard-to-predict, as they have high membership scores on the target model. We refer to such non-members as hard-to-calibrate samples, as they cannot be easily distinguished using calibrated membership scores.
@@
According to Long et al., certain data records are at a higher risk of being exploited by attackers if they have fewer neighbors in the sample space represented by the records available to the attacker [24]. This is because such records may impose a unique influence on the target model. In other words, a member data record with fewer neighbors in the training dataset is more likely to be identified by MIAs. Conversely, data records with more neighbors may lead to more incorrect inferences by MIAs. For example, if a non-member data record has many neighbors in the training dataset of the target model, it is more likely to be incorrectly identified as a member compared to another non-member data record that has fewer neighbors. Therefore, it is speculated that the hard-to-calibrate non-members have more neighbors in the target model’s training dataset. In other words, when comparing two data records with high calibrated membership scores, the one with fewer neighbors is more likely to be a member. Hence, LDC-MIA includes neighbor information to differentiate hard-to-predict members and hard-to-calibrate non-members in the upper right region in Figure 3.
The goal is to lower the calibrated membership score for hard-to-calibrate non-members. A similar method proposed by Long et al. is adopted to calculate neighborhood information. First, a data record is fed into the reference model and the output of the last layer before the softmax layer is collected as its neighborhood vector. Then, cosine similarity is computed between this neighborhood vector and those of all the other data records in the auxiliary dataset. Data records with cosine similarity exceeding a certain threshold are considered neighbors. Finally, the neighborhood information of a target data record $x$ is computed as follows:
@@
, where $\mathbf{v}_{x}$ is the neighborhood vector of the target data record, $\mathbf{v}_{aux_{i}}$ is that of the data records in the auxiliary dataset, $n$ is the size of the auxiliary dataset, and $\theta$ is the similarity threshold value to determine neighbors. We use $\theta=0$ in our attack. Through our experiments in Section 4, we verify that setting $\theta$ to 0 works well for most datasets. The intuition behind this is that the value of cosine similarity is greater than 0 when two data records are positively related. We also know that a model’s inference on a target sample is influenced by its neighbor data records if they exist in the training dataset. Therefore, by using a threshold of 0 for the cosine similarity value, we can better distinguish neighbor and non-related data records. Then, we enhance the membership scores proposed by Watson et al. [37] with neighborhood information as follows:
@@
, where $h$ is the target model, and $g$ is the reference model. This enhancement helps distinguish the hard-to-predict members and hard-to-calibrate non-members by shifting the membership scores differently.
@@
The MIA follows Watson et al.’s approach of using a threshold on calibrated membership scores to determine membership. Based on the figure, it can be observed that the new score can better distinguish members from hard-to-calibrate non-members manifested in improved TPRs at the same low FPRs. It is important to note that we only compared the TPR in the low FPR region because this is the region where MIAs are considered practical. The comparison results suggest that neighborhood information is a valuable and useful component in membership scores.
Different MIA score thresholds are needed for accurately classifying samples with different labels. One of the main objectives of an attacker is to achieve high precision in identifying member data records in the training dataset. To achieve this goal, the attacker strives to differentiate between members and non-members as much as possible. Many existing MIAs rely on a threshold value of the membership scores to distinguish between members and non-members. However, the divergence of membership scores in a target model is influenced not only by the hardness and neighborhood information of a data record but also by its assigned label. This could be attributed to the different distributions of easy and hard data records across various classes. As a result, we believe that determining the most suitable threshold of membership scores for each class can further enhance the accuracy of MIAs. To illustrate this, we present an example of the membership scores of data records in CIFAR-10 belonging to different classes in Figure 5, using the enhanced calibrated membership score calculated by Eq 8. In this example, we use the VGG-16 as both our target and reference models. Both models are trained until they achieve their highest accuracy values on the test dataset.
@@
Figure 6: Workflow of LDC-MIA.

### v3

@@ Abstract
Machine learning models, in particular deep neural networks, are currently an integral part of various applications, from healthcare to finance. However, using sensitive data to train these models raises concerns about privacy and security. One method that has emerged to verify if the trained models are privacy-preserving is Membership Inference Attacks (MIA), which allows adversaries to determine whether a specific data point was part of a model’s training dataset. While a series of MIAs have been proposed in the literature, only a few can achieve high True Positive Rates (TPR) in the low False Positive Rate (FPR) region ($0.01\%\sim 1\%$). This is a crucial factor to consider for an MIA to be practically useful in real-world settings. In this paper, we present a novel approach to MIA that is aimed at significantly improving TPR at low FPRs. Our method, named learning-based difficulty calibration for MIA (LDC-MIA), characterizes data records by their hardness levels using a neural network classifier to determine membership. The experiment results show that LDC-MIA can improve TPR at low FPR by up to 4x compared to the other difficulty calibration-based MIAs. It also has the highest Area Under ROC curve (AUC) across all datasets. Our method’s cost is comparable with most of the existing MIAs, but is orders of magnitude more efficient than one of the state-of-the-art methods, LiRA, while achieving similar performance.
@@ 1 Introduction
Figure 1: The figure shows data records of Airplane and Cat classes in the CIFAR-10 dataset. Each record is represented by a marker indicating its membership type. The y-axis shows the calibrated membership scores, and the x-axis shows the membership scores on the target model. The membership score is determined as the negative of cross-entropy loss values. The calibrated membership score is the difference between a data record’s membership score on the target model and the reference model. In traditional MIAs, data samples with higher membership scores are more likely to be members, while in difficulty calibration-based MIAs, data samples with higher calibrated membership scores are more likely to be members.
@@
To tackle this issue, Watson et al. proposed using a reference model [37], which is trained on data from the same distribution as the target model’s training dataset. By calculating the difference between the loss values obtained from the target and reference models, the reference model helps calibrate the target model’s behavior on a data record. This approach is an example of difficulty calibration-based attacks, one of the most advanced MIAs. To further improve the attack performance, Carlini et al. designed a Likelihood Ratio Attack (LiRA) [3]. LiRA utilizes multiple shadow models to estimate the distribution of loss values on a target data record for models that are either trained or not trained on this data sample.
Existing difficulty calibration-based MIAs rely on a single metric, such as membership or calibrated membership scores, to differentiate between members and non-members. However, such approaches have certain limitations since single metrics may display variations or anomalies that make it difficult to distinguish members from non-members. Therefore, we propose using multiple metrics and information to analyze from different perspectives simultaneously. Our proposed attack [^1] The implementation of LDC-MIA is available at https://github.com/horanshi/LDC-MIA. achieves this by adopting a learning-based approach that utilizes multiple features. We aim to achieve high TPRs at low FPRs and high AUC while minimizing the cost for attackers.
We consider two types of costs in our attack: the training cost and the data cost. The training cost is mainly dependent on the number of attacking models and their complexities, while the data cost involves the amount of data needed to train them. We strive to minimize both in our design. As mentioned by Carlini et al. [3], MIA can also be used as an auditing tool for ML models. For example, a company may use MIA to examine all the training data to detect privacy leakage before releasing a commercial model. Cost is an important factor in such scenarios.
@@
In our discussions, we classify data records based on their difficulty levels. We refer to data records whose class labels can be easily predicted correctly by both the target and the reference models easy-to-predict samples, while those whose labels are difficult to predict correctly are called hard-to-predict samples. To illustrate this, we provide an example in Figure 1. There are also some hard-to-calibrate samples in this figure, which we will explain in Section 3.2. Based on the figure, we can make a few important observations. First, suppose we only consider the membership scores of the target samples on the target model (as shown on the x-axis). In that case, many non-members overlap with members because they are easy-to-predict samples and thus have low membership scores. The members in this overlap can either be easy-to-predict or hard-to-predict samples. Using the calibrated membership scores (as shown on the y-axis) increases the gap between the easy-to-predict non-members and hard-to-predict members. However, if we only consider the calibrated membership scores, the easy-to-predict members and the hard-to-predict non-members may overlap since both groups may have low calibrated membership scores. Therefore, both membership and calibrated membership scores are useful in distinguishing members from non-members. Second, if we use a fixed threshold value on the calibrated membership scores to differentiate between members and non-members, as Watson et al. did, it may not work well for both classes since they have different optimal cut-off points. This indicates that the hardness levels of data records are not universal across different classes. Therefore, adopting a more intelligent approach to determining the threshold values is necessary. Third, data distribution also plays an important role in determining how difficult it is for a data record to be correctly classified, in addition to its intrinsic characteristics. As revealed by Long et al. [22], the more neighbors a data record has in the training dataset, the easier it is for it to be correctly classified. This also means that the data record is more likely to be determined as a member by attackers, regardless of its membership.
@@
Based on these observations, we propose developing a classifier that can learn to calibrate difficulty based on the membership score on the target model, the calibrated membership score, the label of the target data record, and its neighborhood information. To train this classifier, we can use a shadow target model and a reference model trained with data records that share the same distribution as those belonging to the target model training dataset. The shadow target model would mimic the target model’s behavior in classifying members and non-members. We call the proposed attack LDC-MIA. The main contributions of this paper are threefold. (1) The proposed attack significantly improves the TPR at low FPR while minimizing the cost for attackers. We only require one shadow model and one reference model to improve the TPR. The classifier we build is a simple model with three fully connected layers. (2) We conduct a comprehensive characterization of the data records’ hardness levels and use these characters to train a neural network for determining membership. This learning-based calibration approach can be easily extended to integrate other features without requiring significant retraining efforts. (3) Through extensive evaluations, we provide insights into each character’s contributions to the success of our proposed attack.
@@
We conduct extensive experiments to evaluate the performance of LDC-MIA on various datasets. Specifically, we measure the TPR at low FPRs ranging from $0.01\%$ to $1\%$. This metric helps us evaluate the model’s ability to correctly identify positive instances while minimizing the number of false positive predictions for practical use. Our results show that our proposed attack achieves the highest AUC across all datasets compared to state-of-the-art MIAs and improves TPR up to 4x. In addition, we measure the precision-recall curve to analyze how well the model performs across different recall levels while maintaining high precision. The results indicate that LDC-MIA consistently produces the highest precision values for different recall values across all datasets. For instance, LDC-MIA identifies $52.72\%$ of the members with a precision of $80\%$, significantly higher than other MIAs can achieve.
@@ 3.2 Design intuition
Recent state-of-the-art attacks [3, 37] have explored the difficulty level of each data record and applied parametric calibration to improve the attack performance in the low FPR region. These attacks are similar in design to our proposed method in that each target data record is individually considered when performing attacks. However, these works provided limited discussions on the impact of calibration on different types of data records with varying intrinsic properties. Inspired by this gap, we categorize member and non-member data records into five categories based on their difficulty levels for label predictions and calibrations, as shown in Figure 3: hard-to-predict member/non-member, easy-to-predict member/non-member, and hard-to-calibrate non-member. The definition of easy-to-predict and hard-to-predict samples can be found in section 1. We refer to the non-members that have high calibrated membership scores as hard-to-calibrate samples, as their membership cannot be accurately determined by the existing difficulty calibration based MIAs. This figure is similar to Figure 1, and we have divided it into four regions using a fixed membership score threshold and a fixed calibrated membership score threshold. By analyzing this figure, we can get the following insights.
@@
Even though FPR can be reduced by carefully selecting a threshold for the calibrated scores, it is difficult to eliminate all of them. In Figure 3, it can be seen that many non-members overlap with members along the $y$-axis. Difficulty calibration based MIAs have improved traditional MIAs, specifically the TPR in the low FPR region. This is achieved by identifying hard-to-predict members more accurately by carefully selecting threshold values for calibrated membership scores. However, decreasing the number of false positives remains challenging since many non-members overlap with members, as demonstrated in Figure 3. This can be seen more clearly in Figure 2, which shows the distribution of calibrated membership scores for members and non-members. In this figure, the calibrated membership scores are on the $x$-axis and share the same values with that of Figure 3. The $y$-axis shows the number of members and non-members with corresponding calibrated scores. The figure highlights that many members have low calibrated scores, making them easy-to-predict members. These members overlap with non-members near the line where the calibrated score equals 0. Within this region, the easy-to-predict members overlap with both hard-to-predict and easy-to-predict non-members, as both groups have similar outputs on the target and the reference models.
@@
Two important observations can be made from the figure. Firstly, attackers are likely to encounter easy-to-predict members in real-world attacks. Therefore, identifying these samples can significantly improve TPR. Secondly, existing difficulty calibration based MIAs may fail to isolate such members from non-members by only considering the calibrated membership scores. However, Figure 3 shows that adjusting the thresholds for membership and calibrated membership scores simultaneously can make it easier to differentiate between members and non-members. This suggests that the membership scores on the target model are still useful in addition to the calibrated membership scores. Based on this observation, we utilize membership scores on the target model in LDC-MIA to exclude the hard-to-predict non-members. This not only helps identify hard-to-predict members but also easy-to-predict members, thus improving TPR in all FPR regions.
@@
Figure 3: Based on the calibrated membership scores and membership scores on the target model (VGG-16), we group the target samples from the CIFAR-10 dataset into five categories: hard-to-predict member/non-member, easy-to-predict member/non-member, and hard-to-calibrate non-member.
Neighbor information is also important. In Figure 3, it is clear that there are some hard-to-calibrate non-members. These samples are not easy-to-predict as they have low membership scores on the reference model, nor are they hard-to-predict as they have high membership scores on the target model. The most plausible explanation for this phenomenon is that these non-members have more neighbors in the members than other non-member samples. Neighbors are determined by the similarity between two data samples. Specifically, we can input two data samples into a model and compute the cosine similarity of the output vectors of the last layer before the softmax layer. If their cosine similarity exceeds a certain threshold, then they are considered neighbors.
@@
According to Long et al., certain data records are are more vulnerable to being identified as members by attackers if they have fewer neighbors in the training data [22]. This is because such records may display unique characteristics that the target model can overfit to, making them easier to identify. On the other hand, data records with more neighbors may lead to incorrect membership inferences by MIAs. Therefore, a non-member data record with many neighbors in the target model’s training dataset is more likely to be misidentified as a member than another non-member sample with fewer neighbors. This means that if we have two samples with similarly high membership scores, the one with fewer neighbors is more likely to be a member. Motivated by this, LDC-MIA includes neighbor information to differentiate the hard-to-predict members and hard-to-calibrate non-members in the upper right region in Figure 3.
We aim to lower the calibrated membership score for hard-to-calibrate non-members, and to achieve this, we rescale the calibrated membership scores with neighborhood information. However, there are two challenges we need to address when calculating the neighborhood information. Firstly, attackers do not have access to the training dataset of the target model and thus are unable to compute cosine similarities with data samples that are members. Secondly, attackers have no access to the output vectors of the last layer before the softmax layer in the target model, making it impossible to exploit the target model to compute neighborhood information. To address these challenges, we make two important assumptions. Firstly, we assume that if a data sample has more neighbors in the auxiliary dataset, it will also have more neighbors in the training dataset of the target model. This assumption is reasonable since the auxiliary dataset follows the same distribution as the training dataset. Secondly, we assume that if two data samples are neighbors using the output vectors obtained from the target model, they will also be neighbors using the output vectors obtained from the reference model. This intuitive assumption holds when the target and reference models share the same architecture. Based on these assumptions, we can then leverage the reference model and the auxiliary dataset to approximate the neighborhood information of a target data record $x$:
@@
, where $\mathbf{v}_{x}$ is the output vector of the target data record, $\mathbf{v}_{aux_{i}}$ is that of the data records in the auxiliary dataset, $n$ is the size of the auxiliary dataset, and $\theta$ is the similarity threshold value to determine neighbors. We use $\theta=0$ in our attack. Through our experiments in Section 4, we verify that setting $\theta$ to 0 works well for most datasets. The intuition behind this is that the value of cosine similarity is greater than 0 when two data records are positively related. Then, we enhance the membership scores proposed by Watson et al. [37] with neighborhood information as follows:
@@
, where $h$ is the target model, and $g$ is the reference model.
@@
The MIA follows Watson et al.’s approach of using a threshold on calibrated membership scores to determine membership. Based on the figure, it can be observed that the new score can better distinguish members from hard-to-calibrate non-members manifested in improved TPRs at the same low FPRs. It is important to note that we only compared the TPR in the low FPR region because this is the region where MIAs are considered practical. The comparison results suggest that neighborhood information is a valuable component in membership scores.
Different MIA score thresholds are needed for accurately classifying samples of different labels. One of the main objectives of an attacker is to achieve high precision in identifying member data records in the training dataset. To achieve this goal, the attacker strives to differentiate between members and non-members as much as possible. Many existing MIAs rely on a threshold value of the membership scores to distinguish between members and non-members. However, the divergence of membership scores in a target model is influenced not only by the hardness and neighborhood information of a data record but also by its assigned label.
@@
Figure 6: Workflow of LDC-MIA. During the training phase, $\mathcal{D}^{train}_{shadow}$ and $\mathcal{D}^{heldout}_{shadow}$ serve as members and non-members for the shadow model, respectively. They are input into the shadow target model $h$ and the reference model $g$ to obtain membership scores $s(h,(x,y))$ and $s(g,(x,y))$. The calibrated membership score $s^{\text{cal}}(h,g,(x,y))$ is obtained by using Eq 8. Finally, we use $s(h,(x,y))$, $s^{\text{cal}}(h,g,(x,y))$, and the class labels as features to train the MIA classifier. For $\mathcal{D}^{train}_{shadow}$ and $\mathcal{D}^{heldout}_{shadow}$, their ground truth labels in the classifier training are “in” or “out”, respectively. During the inference phase, we follow the same procedure to obtain the features of the target data sample $(x_{target},y_{target})$ except that the target model is used rather than the shadow target model.
This could be attributed to the different distributions of easy and hard data records across various classes. As a result, we believe that determining the most suitable threshold of membership scores for each class can further enhance the accuracy of MIAs. To illustrate this, we present an example of the membership scores of data records in CIFAR-10 belonging to different classes in Figure 5, using the enhanced calibrated membership score calculated by Eq 8. In this example, we use the VGG-16 as both our target and reference models. Both models are trained until they achieve their highest accuracy values on the test dataset.